Integrate ∫(sinx/cos 2 x) dx

Integrate
∫(sinx/cos2x) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int\ \frac{sinx}{cos^2(x)}dx
cosx = t
-sinxdx = dt
I = \int\ \frac{-1}{t^2}dt
I = \frac{1}{t} + constant
I = \frac{1}{cosx} + constant
I = secx + constant